Shift report 20020113 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Jan 13 2002 - 00:11:20 EST

  • Next message: Apache: "Shift report 20020111 16:00-24:00"

    ------------------------------------------------------------
    Date: Sun Jan 13 00:11:20 2002, Shift: 16:00-24:00
    Supervisor: hiro, with: 
    Log book#12 pages 52 - 52
    
    Shift Summary:
    ------------------------------------------------------------
    1.  Beam for most of the shift.
      At the end of the shift, pol flipper study.
    2.  Run 6432
          MRS 4k A at 45deg
          FS  1/3 B at 4deg
        Finished the settings.
        Since DAQ could not handle the events,
        the scale-down of trigger 6 was increased
        from 1 to 5.
    
    
        Run 6434
          MRS 3k A at 60 Deg
          FS 1/15 B at 20 Deg
    3.  None detected.
      
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Jan 14 2002 - 11:31:10 EST